Dear Richard dear all,

please find below a short summary of the activities today on the mountain related to the AO system. Italian people working at LBTO were:

Arcetri: A. Puglisi, P. Stefanini, E. Pinna, A. Riccardi, R. Briguglio, S. Esposito, L. Busoni, M. Xompero
LBTO: G. Brusa
MG: M. Andrighettoni
ADS: E. Anaclerio

Total # of person: 11

W unit

1) The alignment laser unit has been mounted on the board in place of the input lens. The system misalignment with the laser beam has been measured to be less then 2 arcminutes. This misalignment is inside our specification. So the laser can be taken as reference for the alignment with the rotator axis. The alignment of the board is planned for tomorrow afternoon after installation of IRTC.

2) The installation of WFS board cover has shown that some modification are required to install and unistall it easily in particular if the operation is done when the board is mounted on the telescope. Modification are planned to be done in the next 2 days.


AO control SW

1) The fiber connection between the ASM and the W unit has been tested successfully.


LBT672a

1) measurement of least-clearance point position of the unit LBT672a (crate fittings and ribbon cable clips) to be compared with unit B to prove the equivalence in terms of collision with the hub. The comparison will be performed using equivalent measurements on unit B when the AdSec crew will be back in Arcetri (about 1 mo)
2) the yesterday activity for fixing actuators plus the known presence of ~40um-size contamination reported on 26 Feb required a dismounting of the shell for cleaning. After cleaning, the procedure for checking the gap contamination showed spots lower than 35um that is our threshold for releasing the system
3) cleaning of unit, closing the shroud and mounting the "miniskirt" (lateral protection for reference plate and dust contamination)
4) setting the retention ring to 400um as standard configuration
5) installation of second level cover for dust protection over the shell central membrane: a 125um-thick Mylar cover shaped to avoid axial stiffness and held by central shaft of the membrane interface to the reference body (inside the protection cup for central membrane)
6) installation of protection cup
7) a first slewing test showed a reset of the shell because an incorrect setting of overcurrent safety protection (fixed)
8) just before releasing the unit and the cabinet, an extended functionality check in terms of: communication, elevation slewing test stability and step responses was successfully performed to be sure releasing a functioning system in the telescope configuration.
9) cabinet released at 18:00 for telescope mounting (3h of delay).


Next update tomorrow

All the best

Simone

-- ArmandoRiccardi - 02 Apr 2010
  • Rerotator Coating:
    rerotator-coating.jpg
Topic revision: r1 - 02 Apr 2010, ArmandoRiccardi
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback